OpenStudy (anonymous):

Express 2 yards to 45 feet as a ratio ? ( In simplest form )

OpenStudy (johnweldon1993):

\[\large \frac{2 \text{yards}}{6 \text{feet}} = \frac{x}{45 \text{feet}}\]

OpenStudy (anonymous):

Would The Answer Be \[\frac{ 2 }{ 15 }\]?

OpenStudy (johnweldon1993):

Not quite...when we cross multiply that...we get \[\large 90 = 6x\] \[\large x = 15\] so there are 15 yards in 45 feet

OpenStudy (anonymous):

So Whats It As A Ratio In Simplest Form ? Im Confused . . .

OpenStudy (anonymous):

?